# Quantum Calculations of the Cavity Shift in Electron Magnetic Moment Measurements

### v1

@@ 1 Introduction
Measurements of the electron magnetic moment $\mu\equiv\mu_{B}g/2$ have reached a fractional uncertainty of $1.3\times 10^{-13}$, making it the single most precisely measured property of any fundamental particle PhysRevLett.130.071801 ; Fan:2022oyb . Such measurements directly probe five-loop corrections in quantum electrodynamics PhysRevD.97.036001 ; PhysRevLett.109.111807 ; PhysRevD.100.096004 . Future refinements could improve the precision by over an order of magnitude PhysRevLett.126.070402 ; PhysRevA.103.022824 ; PhysRevA.111.042806 , which would yield better sensitivity to generic physics beyond the Standard Model than that of the leading muon g2g-2 measurements Muong-2:2025xyk .
The leading technique for measuring the electron $g$-factor involves making precise measurements of the energy levels of a single electron in a magnetic field $B$ Brown:1985rh ; Gabrielse:2025jep , as depicted in Fig. 1. The energy eigenstates are two ladders of Landau levels $|n,\downarrow\rangle$ and $|n,\uparrow\rangle$, with the Landau levels and spin-flip transitions ideally separated by the cyclotron and spin precession angular frequencies respectively,
@@
In the early 1980s, several groups attempted such a calculation in toy setups with one or two infinite conducting plates, and found comparable shifts of both $\omega_{c}$ and $\omega_{s}$, in disagreement with the classical calculation and with each other Barton_1977 ; PhysRevD.30.2356 ; PhysRevLett.54.742 ; PhysRevD.34.1429 . By the late 1980s, consensus was reached that these conflicting results were due to errors such as computing gauge-dependent quantities, improperly using hard cutoffs, and failing to account for mass renormalization PhysRevD.32.729 ; Barton_1988 ; PhysRevD.36.2181 ; Kreuzer_1988 . Ultimately, the quantum calculations agreed with the classical result.
Following the infinite-plate calculations, the cavity shift was derived classically for an electron at the center of a cylindrical PhysRevA.32.3204 or spherical PhysRevA.34.2638 cavity. (See also Ref. PhysRevA.83.052122 for a generalization to arbitrary points in a cylindrical cavity, and Ref. kinoshita1990quantum for a pedagogical review.) Here the parallel-plate results are not useful, since these cavity shapes confine the electron to a finite volume, changing the photon mode structure much more dramatically. Furthermore, the classical calculations require a careful subtraction of the electron’s divergent self-field from the classical Green’s function. This was the basis for Kramers’ original attempts at renormalizing electromagnetism Kramers1938 ; schweber1994qed , and in the present context, it can be done exactly for an ideal sphere or cylinder. However, it is unclear how to generalize the subtraction to less ideal or symmetric cavity geometries. Finally, these calculations do not naturally accommodate cavity modes with different quality factors, which is important for future measurements.
@@
In Sec. 3, we describe the electron with nonrelativistic quantum mechanics, and use second-order perturbation theory to compute the energy level shift $\Delta\omega_{c}^{\text{cav}}$, as a sum over cavity modes. We also compute the shift $\Delta\omega_{c}^{\text{free}}$ in free space, as an integral over plane waves. Both these quantities are linearly divergent, and the physical cavity shift is their difference. This derivation is similar in spirit to Bethe’s nonrelativistic calculation of the Lamb shift, though in our case some matrix elements are easier to evaluate analytically. Famously, after performing such subtractions, Bethe showed that the Lamb shift is only logarithmically divergent in the nonrelativistic theory, allowing him to find a decent estimate of its magnitude PhysRev.72.339 . However, finding the precise numerical value required a renormalized calculation in the relativistic theory, including a nontrivial matching of the high-energy and low-energy descriptions of the system at a scale comparable to the electron mass, as reviewed in Refs. weinberg ; schweber1994qed ; itzykson2012quantum .
@@ 2 General Considerations
Many effects can potentially shift $\omega_{c}$ at the $10^{-13}$ level, so we will begin by reviewing and estimating them. We will show that many are actually negligible, or can be accounted for independently of the cavity shift, simplifying the analysis of later sections. We largely follow Ref. Brown:1985rh , but also incorporate more recent theoretical results. Where specific choices of parameters such as magnetic field and cavity size are needed, we choose values relevant to the latest iteration of the experiment PhysRevLett.130.071801 ; Fan:2022oyb .
@@
Since the quadrupole potential also provides a radial force, the physical cyclotron frequency $\omega_{c}$ differs from $\omega_{c}^{0}$ by a fractional amount of order $\omega_{z}^{2}/(\omega_{c}^{0})^{2}\sim 10^{-6}$. However, by measuring $\omega_{c}$, $\omega_{z}$, and $\omega_{m}$ in the presence of the potential, one can infer the value $\omega_{c}^{0}$ that would occur in the absence of the potential using the “invariance theorem” Brown:1985rh
@@
which holds even if the trapping potential is misaligned or otherwise imperfect. As discussed in Ref. Brown:1985rh , $\omega_{c}^{0}$ is essentially the quantity that should be used to determine $g$ via Eq. (1.2). It will turn out that the dominant effect of the cavity shift is through its effect on $\omega_{c}$, but we will note when relevant how shifts of $\omega_{z}$ and $\omega_{m}$ affect the determination of $\omega_{c}^{0}$.
@@
The above discussion treats the electron’s motion as nonrelativistic. The leading relativistic effects on the frequencies of the lowest cyclotron levels, computed in Ref. Brown:1985rh , have a fractional size $\Delta\omega_{c}/\omega_{c}\sim\omega_{c}/m\sim 10^{-9}$. A complementary perspective using classical equations of motion is given in Ref. ketter_2014 .
Intuitively, the main effects at this order are the increased relativistic momentum $\mathbf{p}=\gamma m\mathbf{v}$, which slows down the cyclotron motion, and Thomas precession, which shifts the spin precession frequency. The effects of axial and magnetron motion on the measurement of $g$ are suppressed by multiple powers of $\omega_{z}/\omega_{c}$, and are thus negligible. Relativistic effects at the next order were computed in Ref. Wienczek:2022cwa using the Foldy–Wouthuysen transformation, but these contribute at most $(\omega_{c}/m)^{2}\sim 10^{-18}$ and are also negligible. Thus, we may neglect all relativistic effects besides the few mentioned above, which are simply accounted for by shifting the value of $\omega_{c}$.
Famously, one-loop corrections in QED affect the spin precession frequency at the level $\Delta\omega_{s}/\omega_{s}\sim\alpha/2\pi\sim 10^{-3}$. However, loop corrections generally have a much smaller effect on $\omega_{c}$. First, as discussed in Ref. Jentschura:2023irs , vacuum polarization effects are exponentially suppressed, because they only modify the potential near charges, and the trapped electron is macroscopically far away from other charges. Second, the one-loop correction to the electron self-energy yields an effect analogous to the Lamb shift. Since $\omega_{c}/m$ plays the role of $Z\alpha$ in an atom, one would expect $\Delta\omega_{c}/\omega_{c}\sim\alpha\,(\omega_{c}/m)\log(m/\omega_{c})\sim 10^{-10}$. However, this energy level shift turns out to be independent of the cyclotron energy level to leading order, and thus does not affect measurements of $\omega_{c}$ Jentschura:2023irs ; PhysRev.96.523 ; PhysRevD.8.3446 . The leading state-dependent shift appears only at the next order in the strength of the $B$-field, giving a negligible $\Delta\omega_{c}/\omega_{c}\sim\alpha\,(\omega_{c}/m)^{2}\log(m/\omega_{c})\sim 10^{-19}$.
@@
To address this, the Penning trap is placed inside a cavity, which modifies the free space photon modes to discrete cavity modes. When $\omega_{c}$ coincides with the angular frequency of a cavity mode, the damping rate is resonantly enhanced by the mode quality factor $Q\gtrsim 10^{3}$, and when $\omega_{c}$ is well away from a resonant mode, the contribution of that mode to the decay rate is suppressed by $1/Q$. By tuning $\omega_{c}$ to sit between cavity modes, the decay rate can be sufficiently reduced. The length scale of the cavity is $R\simeq 4\,\mathrm{mm}$; in the cylindrical cavity used for the most recent measurement, $R$ is roughly both the radius and the half-length Fan:2022oyb . This value is chosen so that $\omega_{c}$ is comparable to that of fairly low-lying cavity modes, $\omega_{c}R\sim 10$, where the modes can be unambiguously identified, and their $Q$-factors may be reliably measured.
@@
The cavity suppresses the imaginary part of the electron’s self energy, but by the general logic of the Kramers–Kronig relations, analyticity implies that a shift in the imaginary part is inevitably accompanied by a shift in the real part. Classically, the cyclotron frequency shifts because the radiation fields produced by the electron reflect off the cavity walls and act back on the electron PhysRevD.32.729 . As we will see, there are several parts of this “cavity shift”, though only one is currently relevant.
@@
We emphasize that the longitudinal and transverse cavity shifts are independent effects, which must be summed to yield the total cavity shift. As we will discuss in Sec. 7, the longitudinal cavity shift dominates at low frequencies, and is therefore more important for trapped ion experiments PhysRevA.40.6308 ; vandyck2006231 ; BASE:2014drs ; BASE:2016yuo ; Schneider:2017lff ; PhysRevLett.119.033001 , where it is called the “image charge” shift. In this work, we will focus exclusively on the transverse cavity shift, which arises from the fields of cavity modes. As we will see, this effect is more subtle because the mode sum must be renormalized.
@@
There are several other components of the cavity shift, but they are negligible. First, we have neglected the magnetic field $B_{T}^{\prime}\sim E_{T}^{\prime}$, because the magnetic force is smaller by a factor of $v_{c}\sim\sqrt{\omega_{c}/m}$. Second, the axial and magnetron frequencies are also shifted, by PhysRevA.32.3204 ; Jentschura:2023mbv
@@
which may become relevant in the near future. Finally, the spin precession frequency can be modified by the field of the electron’s image magnetic dipole moment $\mu\sim q/m$. As was first argued in Ref. PhysRevD.32.729 , since the radiative magnetic field is $B_{T}^{\prime}\sim\mu\omega_{s}^{2}/(4\pi R)$, we have
@@
The cavity shift has only been calculated for a closed cavity in cylindrical PhysRevA.32.3204 and spherical PhysRevA.34.2638 geometries. In both cases, $\mathbf{E}_{T}^{\prime}$ was evaluated at the electron’s position using the cavity’s exact Green’s function, with the electron’s divergent self-field subtracted. In the spherical case, a simple analytic form for the Green’s function exists, while for the cylindrical case, it can be written as an infinite series by first considering the Green’s function due to two infinite parallel plates, then adding the contribution due to the cylinder’s curved walls. The derivation relies on the symmetry of the cavity, which allows the use of the method of images. In both cases, it is possible to account for finite quality factors by shifting the frequency of the Fourier-space Green’s function into the complex plane, but one must assign the same quality factor to all modes, or in the case of the cylinder, a uniform $Q$-factor to all TE modes and another $Q$-factor to all TM modes.
In practice, the quality factors are not uniform, and the modes do not take their ideal forms, due to imperfections of the cavity geometry and the effects of the electrodes that generate the trapping potential. Thus, in practice the cavity is characterized by measuring properties of its modes. Given this information, one can compute the cavity shift by considering the classical excitation of each cavity mode $n$, giving a result of the form PhysRevA.37.4163
@@
where $c_{n}$ is a mode-dependent order-one coefficient. One can then estimate the cavity shift by summing over a few modes with $\omega_{n}$ close to $\omega_{c}$, as was suggested in Refs. dehmelt1984 ; dehmelt1992 .
@@ 3.1 Electron and Photon States
The Hamiltonian of a nonrelativistic electron in a Penning trap can be solved exactly Brown:1985rh , and we briefly review the results here. To warm up, first suppose there is only a uniform magnetic field, and work in symmetric gauge, where $\mathbf{A}_{c}=\frac{1}{2}(\mathbf{B}\times\bm{\rho})$ with $\bm{\rho}$ the radial vector in cylindrical coordinates. The classical equations of motion are linear, which implies that the quantum states can be expressed as a combination of harmonic oscillators at angular frequency $\omega_{c}^{0}$. To do this, we define the operator
@@
Including the quadrupole potential, the components of the mechanical momentum may be expanded in cyclotron, magnetron, and axial ladder operators as Brown:1985rh
@@
$\mathbf{u}_{\sigma,0\nu p}(\bm{0})$ $\equiv u_{\sigma,\nu p}^{||}\hat{\mathbf{z}}$ (3.18)

### v2

@@ 1 Introduction
Measurements of the electron magnetic moment $\mu\equiv\mu_{B}g/2$ have reached a fractional uncertainty of $1.3\times 10^{-13}$, making it the single most precisely measured property of any fundamental particle Fan et al. (2023); Fan (2022). Such measurements directly probe five-loop corrections in quantum electrodynamics Aoyama et al. (2018). [^1] This calculation involves diagrams with electron, muon, and tau loops, as well as hadronic and electroweak corrections. There was a $5\sigma$ discrepancy in the calculation of the five-loop QED diagrams without lepton loops Aoyama et al. (2012); Volkov (2019), which was recently resolved Volkov (2024); Aoyama et al. (2025), and found to arise from a bias in the numeric integration procedure. The use of electron g2g-2 to test the Standard Model is currently limited by the $5\sigma$ discrepancy in the leading measurements of $\alpha$ by atom interferometry Parker et al. (2018); Morel et al. (2020). Future refinements could improve the precision by over an order of magnitude Fan and Gabrielse (2021a, b); Fan et al. (2025), which would yield better sensitivity to generic physics beyond the Standard Model than that of the leading muon g2g-2 measurements Aguillard and others (2025).
The leading technique for measuring the electron $g$-factor involves making precise measurements of the energy levels of a single electron in a magnetic field $B$ Brown and Gabrielse (1986); Gabrielse and Venanzoni (2025), as depicted in Fig. 1. The energy eigenstates are two ladders of Landau levels $|n,\downarrow\rangle$ and $|n,\uparrow\rangle$, with the Landau levels and spin-flip transitions ideally separated by the cyclotron and spin precession angular frequencies respectively,
@@
In the early 1980s, several groups attempted such a calculation in toy setups with one or two infinite conducting plates, and found comparable shifts of both $\omega_{c}$ and $\omega_{s}$, in disagreement with the classical calculation and with each other Barton and Grotch (1977); Fischbach and Nakagawa (1984); Svozil (1985); Kreuzer and Svozil (1986). By the late 1980s, consensus was reached that these conflicting results were due to errors such as computing gauge-dependent quantities, improperly using hard cutoffs, and failing to account for mass renormalization Boulware et al. (1985); Barton and Fawcett (1988); Tang (1987); Kreuzer (1988). Ultimately, the quantum calculations agreed with the classical result.
Following the infinite-plate calculations, the cavity shift was derived classically for an electron at the center of a cylindrical Brown et al. (1985) or spherical Brown et al. (1986) cavity. (See also Ref. Hanneke et al. (2011) for a generalization to arbitrary points in a cylindrical cavity, and Ref. Kinoshita (1990) for a pedagogical review.) Here the parallel-plate results are not useful, since these cavity shapes confine the electron to a finite volume, changing the photon mode structure much more dramatically. Furthermore, the classical calculations require a careful subtraction of the electron’s divergent self-field from the classical Green’s function. This was the basis for Kramers’ original attempts at renormalizing electromagnetism Kramers (1938); Schweber (1994), and in the present context, it can be done exactly for an ideal sphere or cylinder. However, it is unclear how to generalize the subtraction to less ideal or symmetric cavity geometries. Finally, these calculations do not naturally accommodate cavity modes with different quality factors, which is important for future measurements.
@@
In Sec. 3, we describe the electron with nonrelativistic quantum mechanics, and use second-order perturbation theory to compute the energy level shift $\Delta\omega_{c}^{\text{cav}}$, as a sum over cavity modes. We also compute the shift $\Delta\omega_{c}^{\text{free}}$ in free space, as an integral over plane waves. Both these quantities are linearly divergent, and the physical cavity shift is their difference. This derivation is similar in spirit to Bethe’s nonrelativistic calculation of the Lamb shift, though in our case some matrix elements are easier to evaluate analytically. Famously, after performing such subtractions, Bethe showed that the Lamb shift is only logarithmically divergent in the nonrelativistic theory, allowing him to find a decent estimate of its magnitude Bethe (1947). However, finding the precise numerical value required a renormalized calculation in the relativistic theory, including a nontrivial matching of the high-energy and low-energy descriptions of the system at a scale comparable to the electron mass, as reviewed in Refs. Weinberg (1995); Schweber (1994); Itzykson and Zuber (2012).
@@ 2 General Considerations
Many effects can potentially shift $\omega_{c}$ at the $10^{-13}$ level, so we will begin by reviewing and estimating them. We will show that many are actually negligible, or can be accounted for independently of the cavity shift, simplifying the analysis of later sections. We largely follow Ref. Brown and Gabrielse (1986), but also incorporate more recent theoretical results. Where specific choices of parameters such as magnetic field and cavity size are needed, we choose values relevant to the latest iteration of the experiment Fan et al. (2023); Fan (2022).
@@
Since the quadrupole potential also provides a radial force, the physical cyclotron frequency $\omega_{c}$ differs from $\omega_{c}^{0}$ by a fractional amount of order $\omega_{z}^{2}/(\omega_{c}^{0})^{2}\sim 10^{-6}$. However, by measuring $\omega_{c}$, $\omega_{z}$, and $\omega_{m}$ in the presence of the potential, one can infer the value $\omega_{c}^{0}$ that would occur in the absence of the potential using the “invariance theorem” Brown and Gabrielse (1986)
@@
which holds even if the trapping potential is misaligned or otherwise imperfect. As discussed in Ref. Brown and Gabrielse (1986), $\omega_{c}^{0}$ is essentially the quantity that should be used to determine $g$ via Eq. (1.2). It will turn out that the dominant effect of the cavity shift is through its effect on $\omega_{c}$, but we will note when relevant how shifts of $\omega_{z}$ and $\omega_{m}$ affect the determination of $\omega_{c}^{0}$.
@@
The above discussion treats the electron’s motion as nonrelativistic. The leading relativistic effects on the frequencies of the lowest cyclotron levels, computed in Ref. Brown and Gabrielse (1986), have a fractional size $\Delta\omega_{c}/\omega_{c}\sim\omega_{c}/m\sim 10^{-9}$. A complementary perspective using classical equations of motion is given in Ref. Ketter et al. (2014).
Intuitively, the main effects at this order are the increased relativistic momentum $\mathbf{p}=\gamma m\mathbf{v}$, which slows down the cyclotron motion, and Thomas precession, which shifts the spin precession frequency. The effects of axial and magnetron motion on the measurement of $g$ are suppressed by multiple powers of $\omega_{z}/\omega_{c}$, and are thus negligible. Relativistic effects at the next order were computed in Ref. Wienczek et al. (2022) using the Foldy–Wouthuysen transformation, but these contribute at most $(\omega_{c}/m)^{2}\sim 10^{-18}$ and are also negligible. Thus, we may neglect all relativistic effects besides the few mentioned above, which are simply accounted for by shifting the value of $\omega_{c}$.
Famously, one-loop corrections in QED affect the spin precession frequency at the level $\Delta\omega_{s}/\omega_{s}\sim\alpha/2\pi\sim 10^{-3}$. However, loop corrections generally have a much smaller effect on $\omega_{c}$. First, as discussed in Ref. Jentschura and Moore (2023), vacuum polarization effects are exponentially suppressed, because they only modify the potential near charges, and the trapped electron is macroscopically far away from other charges. Second, the one-loop correction to the electron self-energy yields an effect analogous to the Lamb shift. Since $\omega_{c}/m$ plays the role of $Z\alpha$ in an atom, one would expect $\Delta\omega_{c}/\omega_{c}\sim\alpha\,(\omega_{c}/m)\log(m/\omega_{c})\sim 10^{-10}$. However, this energy level shift turns out to be independent of the cyclotron energy level to leading order, and thus does not affect measurements of $\omega_{c}$ Jentschura and Moore (2023); Newton (1954); Tsai and Yildiz (1973). The leading state-dependent shift appears only at the next order in the strength of the $B$-field, giving a negligible $\Delta\omega_{c}/\omega_{c}\sim\alpha\,(\omega_{c}/m)^{2}\log(m/\omega_{c})\sim 10^{-19}$.
@@
To address this, the Penning trap is placed inside a cavity, which modifies the free space photon modes to discrete cavity modes. When $\omega_{c}$ coincides with the angular frequency of a cavity mode, the damping rate is resonantly enhanced by the mode quality factor $Q\gtrsim 10^{3}$, and when $\omega_{c}$ is well away from a resonant mode, the contribution of that mode to the decay rate is suppressed by $1/Q$. By tuning $\omega_{c}$ to sit between cavity modes, the decay rate can be sufficiently reduced. The length scale of the cavity is $R\simeq 4\,\mathrm{mm}$; in the cylindrical cavity used for the most recent measurement, $R$ is roughly both the radius and the half-length Fan (2022). This value is chosen so that $\omega_{c}$ is comparable to that of fairly low-lying cavity modes, $\omega_{c}R\sim 10$, where the modes can be unambiguously identified, and their $Q$-factors may be reliably measured.
@@
The cavity suppresses the imaginary part of the electron’s self energy, but by the general logic of the Kramers–Kronig relations, analyticity implies that a shift in the imaginary part is inevitably accompanied by a shift in the real part. Classically, the cyclotron frequency shifts because the radiation fields produced by the electron reflect off the cavity walls and act back on the electron Boulware et al. (1985). As we will see, there are several parts of this “cavity shift”, though only one is currently relevant.
@@
We emphasize that the longitudinal and transverse cavity shifts are independent effects, which must be summed to yield the total cavity shift. As we will discuss in Sec. 7, the longitudinal cavity shift dominates at low frequencies, and is therefore more important for trapped ion experiments Van Dyck et al. (1989, 2006); Mooser et al. (2014); Smorra and others (2017); Schneider and others (2017); Heiße et al. (2017), where it is called the “image charge” shift. In this work, we will focus exclusively on the transverse cavity shift, which arises from the fields of cavity modes. As we will see, this effect is more subtle because the mode sum must be renormalized.
@@
There are several other components of the cavity shift, but they are negligible. First, we have neglected the magnetic field $B_{T}^{\prime}\sim E_{T}^{\prime}$, because the magnetic force is smaller by a factor of $v_{c}\sim\sqrt{\omega_{c}/m}$. Second, the axial and magnetron frequencies are also shifted, by Brown et al. (1985); Jentschura (2023)
@@
which may become relevant in the near future. Finally, the spin precession frequency can be modified by the field of the electron’s image magnetic dipole moment $\mu\sim q/m$. As was first argued in Ref. Boulware et al. (1985), since the radiative magnetic field is $B_{T}^{\prime}\sim\mu\omega_{s}^{2}/(4\pi R)$, we have
@@
The cavity shift has only been calculated for a closed cavity in cylindrical Brown et al. (1985) and spherical Brown et al. (1986) geometries. In both cases, $\mathbf{E}_{T}^{\prime}$ was evaluated at the electron’s position using the cavity’s exact Green’s function, with the electron’s divergent self-field subtracted. In the spherical case, a simple analytic form for the Green’s function exists, while for the cylindrical case, it can be written as an infinite series by first considering the Green’s function due to two infinite parallel plates, then adding the contribution due to the cylinder’s curved walls. The derivation relies on the symmetry of the cavity, which allows the use of the method of images. In both cases, it is possible to account for finite quality factors by shifting the frequency of the Fourier-space Green’s function into the complex plane, but one must assign the same quality factor to all modes, or in the case of the cylinder, a uniform $Q$-factor to all TE modes and another $Q$-factor to all TM modes.
In practice, the quality factors are not uniform, and the modes do not take their ideal forms, due to imperfections of the cavity geometry and the effects of the electrodes that generate the trapping potential. Thus, in practice the cavity is characterized by measuring properties of its modes. Given this information, one can compute the cavity shift by considering the classical excitation of each cavity mode $n$, giving a result of the form Brown et al. (1988)
@@
where $c_{n}$ is a mode-dependent order-one coefficient. One can then estimate the cavity shift by summing over a few modes with $\omega_{n}$ close to $\omega_{c}$, as was suggested in Refs. Dehmelt (1984); Dehmelt et al. (1992).
@@ 3.1 Electron and Photon States
The Hamiltonian of a nonrelativistic electron in a Penning trap can be solved exactly Brown and Gabrielse (1986), and we briefly review the results here. To warm up, first suppose there is only a uniform magnetic field, and work in symmetric gauge, where $\mathbf{A}_{c}=\frac{1}{2}(\mathbf{B}\times\bm{\rho})$ with $\bm{\rho}$ the radial vector in cylindrical coordinates. The classical equations of motion are linear, which implies that the quantum states can be expressed as a combination of harmonic oscillators at angular frequency $\omega_{c}^{0}$. To do this, we define the operator
@@
Including the quadrupole potential, the components of the mechanical momentum may be expanded in cyclotron, magnetron, and axial ladder operators as Brown and Gabrielse (1986)
@@
$\mathbf{u}_{\sigma,0\nu p}(\bm{0})$ $\equiv u_{\sigma,\nu p}^{||}\hat{\mathbf{z}}$, (3.18)